Graph the following conic section
x=y^2+6y+5

Answer by stanbon(75887)

Graph the following conic section
x=y^2+6y+5
----
Not a function, so cannot be shown on this site.
-----
Parabola opening to the right from its vertex.
Vertex occurs at y = -b/(2a) = -6/(2) = -3
x = 9-18+5 = -4
Vertex(-4,-3)
x-intercept:: (5,0)
-----------------------
Cheers,
Stan H.
-------------------